OpenStudy (anonymous):

The Nutty Shoppe has 10 kg of mixed pecans and cashews which sell for $8.40 per kilogram. Cashews alone sell for $8 per kilogram, and pecans sell for $9 per kilogram. How many kilograms of pecans are there in the mix? @The Nutty Shoppe has 10 kg of mixed pecans and cashews which sell for $8.40 per kilogram. Cashews alone sell for $8 per kilogram, and pecans sell for $9 per kilogram. How many kilograms of pecans are there in the mix?@ßαbчღ

OpenStudy (anonymous):

C kg Cashews, P kg pecans C + P = 10 (solving for C is C=10-P) $8.40/kg * 10 kg = $84 8 * C + 9 * P = 84 8 * (10 - P) + 9 * P = 84 (substituting in C=10-P) 80 - 8 * P + 9 * P = 84 (rearrange) P = 84 - 80 = 4 kg pecans and therefore 6 kg of cashews check: 8 * C + 9 * 4 = 84 8 * C + 36 = 84 8 * C = 48 C = 6